By how much should the pressure on a litre of water be changed to compress it by 0.10% ?


1
Expert's answer
2021-09-12T19:00:41-0400
"\Delta p=B\frac{\Delta V}{V}\\
\Delta p=2.2\cdot10^9\frac{0.1}{100}=2.2\cdot10^6\ Pa"